# Scaling Up Active Testing to Large Language Models

### v1

@@ 6 Related Work
The sampling-based approach to active testing that our approach builds on was proposed by Kossen et al, [16], using risk estimators from Farquhar et al, [8]. Earlier work on active testing proposed alternative methods. Bennett & Carvalho, [2], Ji et al, [14], Katariya et al, [15] and Kumar & Raj, [19] explored stratification-based techniques, such as splitting the pool set into strata by model confidence and sampling uniformly within each stratum. Sawade et al, [30] and Yilmaz et al, [37] proposed importance-sampling and Poisson-sampling methods related to that of Kossen et al, [16] but without the latter’s innovations of adaptive, surrogate-based data acquisition and risk estimation that accounts for sampling without replacement. Nguyen et al, [24] studied a special case of active testing focused on human vetting of noisy labels.
Recent years have seen various extensions of this earlier active-testing work. Huang et al, [12] introduced a clustering-based approach that relies on computing predictions through the target model on the pool set, which we have argued is a barrier to scaling up. Su et al, [32] proposed a technique tailored to “dense” recognition tasks in computer vision (e.g., image segmentation and object detection). Yu et al, [38] incorporated active testing into the process of training a model. Ashury-Tahan et al, [1] and Hara et al, [10] proposed methods for model selection.
The idea of reducing model-evaluation costs by using carefully selected test datasets has been highlighted a number of times in recent work. Maynez et al, [23] found that the datasets they studied could be substantially reduced in size, through uniform-random subsampling, while maintaining stable rankings of the models they were comparing. Polo et al, [28], Saranathan et al, [29] and Vivek et al, [35] demonstrated more sophisticated methods for achieving the same goal. The goal in these studies was to take an existing test dataset (with labels for all inputs) and reduce it in a way that supports comparisons between different models. This prior work is therefore complementary to our contribution: we primarily study the problem of acquiring new data with which to evaluate a model of interest; and we use a method that tailors data acquisition to that given model.
@@ 7 Conclusion
We have argued that a significant barrier to principled active testing of large models is the computational cost of deciding which labels to acquire. To address this we have identified key contributors to the cost—training the surrogate model, and computing predictions with the surrogate and target models—and have proposed straightforward cost-saving measures. We have shown that these measures are compatible with effective active testing, producing low-error risk estimates for large language models. On top of this we have demonstrated an estimator of risk-estimation accuracy that can provide an on-the-fly indication of how well active testing is working on a given practical problem. Overall we believe this represents substantial progress towards more efficient evaluation of LLMs, with significant empirical improvements over current standard practice demonstrated.
Appendix A Experiment Details
A.1 Datasets
@@
Our experiments are based on four datasets: Stanford Sentiment Treebank 2 (SST2) ([31], unknown license), Subjectivity (Subj) ([26], Creative Commons Attribution 4.0 International) Financial Phrase-bank (FPB) ([22], Creative Commons Attribution Non Commercial Share Alike 3.0 Unported License), Hate Speech (HS) ([6], Creative Commons Attribution-Share Alike 3.0 Spain License) and Massive Multitask Language Understanding (MMLU) ([11], MIT license).
@@
Each method, including uniform sampling, samples from the same active set, although its size may vary between experiments. We compare error estimates with the error of a separate test dataset, except for the FPB dataset for which we use the error of the active set due to its reduced size. Table˜A.1 summarizes sizes of the active set and test set for each dataset.
Table A.1: Active set and test set sizes for each dataset. Sets are randomly sampled so that they are disjoints.
@@
The Subjectivity dataset is composed of 55k subjective sentences from movie reviews and 55k objective sentences from plot summaries. This dataset is used to determine sentiment polarity. The task is to classify each sentence as either ‘subjective’ or ‘objective’. This dataset is balanced, with $49.5\%$ of objective and $50.5\%$ of subjective sentences. We randomly split the 88k-sentence train set into a 66k-active set and a 44k-test set. Examples for in-context-learning are selected from the separate 22k-sentence test set.
@@
A.2 Models
We employ LLMs from the Llama-2 family ([34], Llama 2 Community License). These are the 7B and 70B versions. We use 8 bit-quantisation for the Llama-2-7B model and half-precision floating-point numbers for the Llama-2-70B model. Kossen et al, [18] showed that these approximations do not significantly affect performance. We additionally employ Gemma-3 4B ([33], Gemma License).
A.3 Prompt formatting

### v2

@@ 6 Related work
The sampling-based approach to active testing that our approach builds on was proposed by kossen2021active, using risk estimators from farquhar2021statistical. Earlier work on active testing proposed alternative methods. bennett2010online, ji2021active, katariya2012active and kumar2018classifier explored stratification-based techniques, such as splitting the pool set into strata (based on a measure of model confidence) and sampling uniformly within each stratum. sawade2010active and yilmaz2021sample proposed importance-sampling and Poisson-sampling methods related to that of kossen2021active but without the latter’s innovations of adaptive, surrogate-based data acquisition and risk estimation that accounts for sampling without replacement. nguyen2018active studied a special case of active testing focused on human vetting of noisy labels.
Recent years have seen various extensions of this earlier active-testing work. huang2025active introduced a clustering-based approach that relies on making predictions with the target model on the pool set, which we have argued is a barrier to scaling up. su2024metaat proposed a technique tailored to “dense” recognition tasks in computer vision (e.g., image segmentation and object detection). yu2024actively incorporated active testing into the process of training a model. ashurytahan2024label and hara2024active proposed methods for model selection.
The idea of reducing model-evaluation costs by using carefully selected test datasets has been highlighted a number of times in recent work. maynez2023benchmarking found that the datasets they studied could be substantially reduced in size, through uniform-random subsampling, while maintaining stable rankings of the models they were comparing. polo2024tinybenchmarks, saranathan2024dele and vivek2024anchor demonstrated more sophisticated methods for achieving the same goal. The goal in these studies was to take an existing test dataset (with labels for all inputs) and reduce it in a way that supports comparisons between different models. This prior work is therefore complementary to our contribution: we primarily study the problem of acquiring new data with which to evaluate a model of interest; and we use a method that tailors data acquisition to that given model.
@@ 7 Conclusion
We have argued that a significant barrier to principled active testing of large models is the computational cost of deciding which labels to acquire. To address this we have identified key contributors to the cost—training the surrogate model, and making predictions with the surrogate and target models—and have proposed straightforward cost-saving measures. We have shown that these measures are compatible with effective active testing, producing low-error risk estimates for large language models. On top of this we have demonstrated an estimator of risk-estimation accuracy that can provide an on-the-fly indication of how well active testing is working on a given practical problem. Overall we believe this represents substantial progress in more efficiently evaluating LLMs.
Acknowledgements
Freddie Bickford Smith was supported by the EPSRC CDT in Autonomous Intelligent Machines and Systems (EP/L015897/1). Tom Rainforth was supported by EPSRC grant EP/Y037200/1.
NeurIPS paper checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We support the claims in Sections 3, 4 and 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We highlight failure cases in Section 5, particularly in Section 5.8.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not present new theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide pseudocode in Algorithms 1 and 2, implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide pseudocode in Algorithms 1 and 2, implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing. All datasets and models used in our experiments are open-source.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify data splits, hyperparameters and other implementation details in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report error bars and coverage probabilities in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide information about computational resources in Section B.7.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics and confirm that this research adheres to the code in all respects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss this in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We credit the creators of the datasets and models we use in Sections 5 and B.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Impact statement
This paper presents work whose goal is to generally advance label- and compute-efficient model evaluation. As this is not limited to a particular area of application, it is difficult to evaluate the impact of this work. Generally speaking, we believe our work could play a role in reducing the computational costs, and thus the CO2 footprint, associated with model evaluation. As is often the case with research, we would caution against careless use of our method in real-world applications.
Appendix B Experiment details
B.1 Datasets
@@
Our experiments are based on five datasets: Stanford Sentiment Treebank 2 (SST2; socher2013recursive; unknown license), Subjectivity (Subj; pang2004sentimental; Creative Commons Attribution 4.0 International License), Financial Phrase-bank (FPB; malo2013good; Creative Commons Attribution Non Commercial Share Alike 3.0 Unported License), Hate Speech (HS; degibert2018hate; Creative Commons Attribution-Share Alike 3.0 Spain License) and Massive Multitask Language Understanding (MMLU; hendrycks2021measuring; MIT license).
@@
Each method, including uniform-random testing, acquires data from the same pool set (a set of candidate data for acquiring, with labels hidden until they are acquired), although its size may vary between experiments. We compare risk estimates against a ground-truth risk, which is computed on a separate test dataset in all cases except for FPB, where we use the pool set due to the small size of the original dataset. Table 1 summarizes sizes of the pool and test sets for each dataset.
@@
Table 1: Pool-set and test-set sizes for each dataset. Sets are randomly sampled so that they are disjoint.
@@
The Stanford Sentiment Treebank 2 dataset consists of 69,000 human-annotated sentences extracted from movie reviews. Each sentence is labeled as either “positive” or “negative”, providing a benchmark for assessing sentiment-analysis capabilities. We use the predefined split of the dataset, concentrating on the training set, which contains 56% positive and 44% negative labels. We randomly select two subsets of 10,000 sentences each from the training set to form the pool and test sets. The instruction used for this dataset is
@@
The string "Answer" is replaced by "Label" for this dataset only.
@@
The Subjectivity dataset is composed of 5,000 subjective sentences from movie reviews and 5,000 objective sentences from plot summaries. This dataset is used to determine sentiment polarity. The task is to classify each sentence as either “subjective” or “objective”. This dataset is balanced, with 49.5% of objective and 50.5% of subjective sentences. We randomly split the 10,000-sentence train set into a 6,000-sentence pool set and a 4,000-sentence test set. Examples for in-context-learning are selected from the separate 2,000-sentence test set. The instruction used for this dataset is
@@
The Financial Phrase-bank dataset consists of 4,800 English financial news articles that were classified as “positive”, “neutral” and “negative” by human experts. We select the training subset of 2,200 sentences for which all 16 annotators agreed. This set contains 13% of negative, 25% of positive and 61% of neutral sentences. Due to its reduced size, the whole set is used as the pool set, and the approximation of the exact loss is the loss over the whole set. The instruction used for this dataset is
@@
The Hate Speech dataset is composed of 9,900 sentences extracted from Stormfront, a white-supremacist forum, from which we select the 9,600 sentences which are classified as “hate” or “no hate” to obtain a binary-classification task. The dataset is highly unbalanced, with 11% classified as hate speech and 89% as non-hate speech. We randomly split this dataset into a pool set of size $\sim$6,000 and a test set of size $\sim$4,000. The instruction used for this dataset is
@@
The Massive Multitask Language Understanding dataset is composed of 116,000 sentences across 57 tasks asssing world knowledge and problem-solving. Each question is multiple-choice, with four choices available. We randomly split this dataset into a pool set of size 7,000 and a test set of size 7,000. The instruction used for this dataset is
@@
B.2 Models
We use the 7B and 70B models from the Llama 2 family (touvron2023llama; Llama 2 Community License). We use 8 bit-quantisation for the 7B model and half-precision floating-point numbers for the 70B model; kossen2024context showed that these approximations do not significantly affect performance. We additionally use Gemma-3 4B (kamath2025gemma; Gemma License).
B.3 Prompt formatting
To evaluate a model on a classification task, we follow the formatting guidelines from kossen2024context in how we generate the input sentence. We begin with a dataset-specific instruction, such as